I am trying to solve the following equation for y:
3y - x
------ = w
y
I don't understand how to isolate the variable "y" because it is in both the numerator and the denominator.

Answer by checkley77(12844)

3y - x
------ = w CROSS MULTIPLY
y
3Y-X=WY SUBTRACT WY & ADD X.
3Y-WY=X FACTOR Y FROM THE LEFT SIDE
Y(3-W)=X DIVIDE BY (3-W)
Y=X/(3-W) ANS.